I do not understand how to put regular square root problems into the simpelest radical form. For example, if i have the Square root of 125 how do i break that down? If you could give me a step by step guide that would be so helpfull for not only this problem but others like it. Thank you.
Found 2 solutions by Alan3354, checkley77:
Answer by Alan3354(69443)

I do not understand how to put regular square root problems into the simpelest radical form. For example, if i have the Square root of 125 how do i break that down? If you could give me a step by step guide that would be so helpfull for not only this problem but others like it. Thank you.
-----------------
Always factor what's under the radical. 125 = 5*5*5
So sqrt(125) = sqrt(5*5*5)
= sqrt((5^2)*5)
= sqrt(5^2)*sqrt(5)
= 5sqrt(5)
--------------
Factor it, look for even exponents, split it up and simplify it.

Answer by checkley77(12844)

sqrt(125)=sqrt(25*5)=sqrt(5*5*5)=5sqrt(5) ans.